# Uncertainty Quantification for LLM-Based Survey Simulations

### v1

@@ B.1 Failure of Exact CLT-Based Intervals under Distribution Shift
$\mathbb{P}\Big{(}\mu\in\mathcal{I}^{\mathsf{syn}}(\widehat{k})\Bigm{|}\widehat{k}\Big{)}\geq 1-\alpha-\sqrt{\frac{2\log(1/\delta)}{m}}.$ (B.1)
@@
$\geq 1-\alpha-\sqrt{\frac{2}{m}}.$
@@
$\bar{k}=\inf\left\{k\in[K]:\mathbb{P}\Big{(}\mu\not\in\mathcal{I}^{\mathsf{syn}}(k)\Big{)}>\alpha+\varepsilon\right\}.$ (B.2)
If $\bar{k}=\inf\emptyset$ does not exist, then there is nothing to prove. Now suppose that $\bar{k}\in[K]$ exists. We will prove that with probability at least $1-\delta$, it holds that $G(\bar{k})>\alpha/2$. When this event happens, we have $\widehat{k}<\bar{k}$, which implies $\mathbb{P}\big{(}\mu\not\in\mathcal{I}^{\mathsf{syn}}(\widehat{k})\bigm{|}\widehat{k}\big{)}\leq\alpha+\varepsilon$ and thus (B.1), thanks to the independence of $\widehat{k}$ and $(\psi,\mathcal{D}^{\mathsf{syn}})$.
@@
$=\mathbb{E}\left[\frac{1}{2}\cdot\mathbb{P}\Big{(}\mu_{j}<\min\mathcal{I}^{\mathsf{syn}}_{j}(k)\Bigm{|}\psi_{j}\Big{)}\right]$
$=\frac{1}{2}\mathbb{P}\Big{(}\mu_{j}<\min\mathcal{I}^{\mathsf{syn}}_{j}(k)\Big{)}.$
@@
$\mathbb{P}\Big{(}\bar{y}_{j}\geq\mu_{j}\text{ and }\mu_{j}>\max\mathcal{I}^{\mathsf{syn}}_{j}(k)\Big{)}=\frac{1}{2}\mathbb{P}\Big{(}\mu_{j}>\max\mathcal{I}^{\mathsf{syn}}_{j}(k)\Big{)}.$
@@
$\mathbb{P}\Big{(}\bar{y}_{j}\not\in\mathcal{I}^{\mathsf{syn}}_{j}(k)\Big{)}$ $\geq\frac{1}{2}\left[\mathbb{P}\Big{(}\mu_{j}<\min\mathcal{I}^{\mathsf{syn}}_{j}(k)\Big{)}+\mathbb{P}\Big{(}\mu_{j}>\max\mathcal{I}^{\mathsf{syn}}_{j}(k)\Big{)}\right]$
$=\frac{1}{2}\mathbb{P}\Big{(}\mu_{j}\not\in\mathcal{I}^{\mathsf{syn}}_{j}(k)\Big{)}=\frac{1}{2}\mathbb{P}\Big{(}\mu\not\in\mathcal{I}^{\mathsf{syn}}(k)\Big{)},$ (B.4)
where the last equality is due to Assumption 2.1. When the event in (B.3) happens,
@@
$\geq\frac{1}{2}\mathbb{P}\Big{(}\mu\not\in\mathcal{I}^{\mathsf{syn}}(\bar{k})\Big{)}-\sqrt{\frac{\log(1/\delta)}{2m}}$ (by (B.4))
@@
$G(k)\leq\frac{1}{m}\sum_{j=1}^{m}\mathbb{P}\Big{(}\bar{y}_{j}\not\in\mathcal{I}^{\mathsf{syn}}_{j}(k)\Big{)}+\sqrt{\frac{\log(n/\delta)}{2m}},\quad\forall k\leq\min\bigg{\{}n,K,~{}\bigg{(}\frac{C}{5\Delta}\bigg{)}^{2}\bigg{\}}$ (B.5)
@@
$\mathbb{P}\Big{(}y^{\mathsf{syn}}_{j}\not\in\mathcal{I}^{\mathsf{syn}}_{j}(k)\Big{)}=\mathbb{P}\left(|y^{\mathsf{syn}}_{j}-\bar{y}^{\mathsf{syn}}_{j,k}|>\frac{C}{\sqrt{k}}\right).$
Since $\bar{y}^{\mathsf{syn}}_{j,k}\sim N(\mu^{\mathsf{syn}}_{j},1/k)$ and $\bar{y}_{j}\sim N(\mu_{j},1/n)$, then
$y^{\mathsf{syn}}_{j}-\bar{y}^{\mathsf{syn}}_{j,k}\sim N\left(\mu_{j}-\mu^{\mathsf{syn}}_{j},~{}\frac{1}{k}+\frac{1}{n}\right).$
@@
$\mathbb{P}\Big{(}\bar{y}_{j}\not\in\mathcal{I}^{\mathsf{syn}}_{j}(k)\Big{)}$ $\leq\mathbb{P}\left(|(\bar{y}_{j}-\bar{y}^{\mathsf{syn}}_{j,k})-(\mu_{j}-\mu^{\mathsf{syn}}_{j})|+\varepsilon>\frac{C}{\sqrt{k}}\right)$
$=2\Phi\left(-\frac{C/\sqrt{k}-\varepsilon}{\sqrt{k^{-1}+n^{-1}}}\right)\leq 2\Phi\bigg{(}-\frac{4C/(5\sqrt{k})}{\sqrt{k^{-1}+k^{-1}}}\bigg{)}$ ($\Delta\leq C/(5\sqrt{k})$)
@@
When $m>\log(n/\delta)/(2\xi^{2})$, substituting (B.6) into (B.5) yields that for all $k\leq\min\{n,K,~{}(\frac{C}{5\Delta})^{2}\}$,
@@
When this happens, we have $\widehat{k}\geq\min\{n,K,~{}(\frac{C}{5\Delta})^{2}\}$.
@@
$=\gamma\cdot\mathbb{P}\Big{(}\theta(\psi)\not\in{\mathcal{S}}^{\mathsf{syn}}(k)\Big{)},$ (B.10)
where the last equality is due to Assumption 2.1. Therefore, when the event in (B.9) happens,
@@ C.1 The OpinionQA Dataset
The original dataset is categorized into topics such as health, crime/security, and political issues. Ideally, we would want to consider questions from the same category to ensure that they are similar enough. However, the category with the most questions has fewer than 200 questions, which is not suited for demonstration. We thus consider all questions. In total, the dataset has 1,442 survey questions, which is too large for our computational resources. We thus selected a subset of questions. First, while the number of choices ranges from 2 to 19, most questions have 5 choices. To give a fair comparison and for simplicity, we therefore only consider questions with 5 choices. Second, not all questions have choices that can be clearly ordered in sentiments; the following question is an example:
@@
In contrast, all example questions in the next section have choices that can be clearly ordered in sentiments. To streamline the process, we ask GPT-4o to determine if a question’s choices can be ordered in sentiments and we keep those that have GPT-4o’s affirmative answer. This leaves us with 546 questions. To compensate for the loss of similarity by pooling questions across various topics and to further reduce our computational cost, we selected 400 questions that are “most similar” to each other by embedding the question statements using OpenAI’s text-embedding-3-small, calculating the mean, and selecting the 400 questions with the smallest Euclidean distance to the mean. Out of these 400 questions, 15 questions have various issues with their choices by manual inspection, so we exclude them. This leaves us with 385 questions. All these questions happen to have at least 400 responses.
@@
Table 1: Categories of surveyees’ features in the OpinionQA dataset.
@@
While LLMs generally provide explanations for the math questions in EEDI dataset, they usually directly provide answers for the OpinionQA dataset; e.g., ‘[[2]]’.
@@ C.3 Visualization of Simulation Sample Size Selection
Figure 5: Visualization of simulation sample size selection for the OpinionQA dataset. Here $\alpha=0.1$, $\gamma=0.5$, $K=100$, $c=\sqrt{2}$ and $M=2$. The red dotted horizontal line is threshold $\gamma\alpha=0.05$. The blue line is $\{L(k)\}_{k=1}^{K}$. Our approach chooses $\widehat{k}$ as the point at which the blue line first up-crosses the red threshold: $\widehat{k}=57$.
@@ D.1 Miscoverage Probability Proxy
Table 5: Means and standard errors of $\widetilde{L}(\widehat{k})$ over 100 random splits for various LLMs and $\alpha$ over the OpinionQA dataset
@@
Table 6: Means and standard errors of $\widetilde{L}(\widehat{k})$ over 100 random splits for various LLMs and $\alpha$ over the EEDI dataset
@@ D.3 Half-Width of Confidence Interval
In Figure 6, Table 9 and Table 10, we present the detailed experiment results for the half-width of the synthetic confidence interval ${\mathcal{S}}^{\mathsf{syn}}(\widehat{k})$.
Figure 6: Half-widths of confidence intervals for various LLMs and $\alpha$’s over the OpinionQA (left) and EEDI (right) datasets.
@@
Table 9: Average half-width (with 95% margin of error in parentheses) of the synthetic confidence interval for various LLMs and $\alpha$ over the OpinionQA dataset. GPT-4o has the shortest confidence interval.
@@
Table 10: Average half-width (with 95% margin of error in parentheses) of the synthetic confidence interval for various LLMs and $\alpha$ over the EEDI dataset. DeepSeek-V3 has the shortest confidence interval.

### v3

@@ B.1 Failure of Exact CLT-Based Intervals under Distribution Shift
$\mathbb{P}\Big{(}\mu\in\mathcal{I}^{\mathsf{syn}}(\widehat{k})\Bigm{|}\widehat{k}\Big{)}\geq 1-(1-2\eta)^{-1}\left(\alpha+\sqrt{\frac{2\log(1/\delta)}{m}}\right).$ (B.1)
@@
$=1-\frac{\alpha}{1-2\eta}-\sqrt{\frac{\pi}{2}}\cdot(1-2\eta)^{-1}\sqrt{\frac{1}{m}}$
$\geq 1-(1-2\eta)^{-1}\bigg{(}\alpha+\sqrt{\frac{2}{m}}\bigg{)}.$
@@
$\bar{k}=\inf\left\{k\in[K]:\mathbb{P}\Big{(}\mu\not\in\mathcal{I}^{\mathsf{syn}}(k)\Big{)}>(1-2\eta)^{-1}\left(\alpha+\varepsilon\right)\right\}.$ (B.2)
If $\bar{k}=\inf\emptyset$ does not exist, then there is nothing to prove. Now suppose that $\bar{k}\in[K]$ exists. We will prove that with probability at least $1-\delta$, it holds that $G(\bar{k})>\alpha/2$. When this event happens, we have $\widehat{k}<\bar{k}$, which implies $\mathbb{P}\big{(}\mu\not\in\mathcal{I}^{\mathsf{syn}}(\widehat{k})\bigm{|}\widehat{k}\big{)}\leq(1-2\eta)^{-1}(\alpha+\varepsilon)$ and thus (B.1), thanks to the independence of $\widehat{k}$ and $(\psi,\mathcal{D}^{\mathsf{syn}})$.
@@
$\geq\mathbb{E}\left[\left(\frac{1}{2}-\eta\right)\cdot\mathbb{P}\Big{(}\mu_{j}<\min\mathcal{I}^{\mathsf{syn}}_{j}(k)\Bigm{|}\psi_{j}\Big{)}\right]$
$=\left(\frac{1}{2}-\eta\right)\cdot\mathbb{P}\Big{(}\mu_{j}<\min\mathcal{I}^{\mathsf{syn}}_{j}(k)\Big{)}.$
@@
$\mathbb{P}\Big{(}\bar{y}_{j}\geq\mu_{j}\text{ and }\mu_{j}>\max\mathcal{I}^{\mathsf{syn}}_{j}(k)\Big{)}\geq\left(\frac{1}{2}-\eta\right)\cdot\mathbb{P}\Big{(}\mu_{j}>\max\mathcal{I}^{\mathsf{syn}}_{j}(k)\Big{)}.$
@@
$\mathbb{P}\Big{(}\bar{y}_{j}\not\in\mathcal{I}^{\mathsf{syn}}_{j}(k)\Big{)}$ $\geq\left(\frac{1}{2}-\eta\right)\cdot\left[\mathbb{P}\Big{(}\mu_{j}<\min\mathcal{I}^{\mathsf{syn}}_{j}(k)\Big{)}+\mathbb{P}\Big{(}\mu_{j}>\max\mathcal{I}^{\mathsf{syn}}_{j}(k)\Big{)}\right]$
$=\left(\frac{1}{2}-\eta\right)\cdot\mathbb{P}\Big{(}\mu_{j}\not\in\mathcal{I}^{\mathsf{syn}}_{j}(k)\Big{)}$
$=\left(\frac{1}{2}-\eta\right)\cdot\mathbb{P}\Big{(}\mu\not\in\mathcal{I}^{\mathsf{syn}}(k)\Big{)}.$ (B.4)
When the event in (B.3) happens,
@@
$\geq\left(\frac{1}{2}-\eta\right)\cdot\mathbb{P}\Big{(}\mu\not\in\mathcal{I}^{\mathsf{syn}}(\bar{k})\Big{)}-\sqrt{\frac{\log(1/\delta)}{2m}}$ (by (B.4))
@@
$G(k)\leq\frac{1}{m}\sum_{j=1}^{m}\mathbb{P}\Big{(}\bar{y}_{j}\not\in\mathcal{I}^{\mathsf{syn}}_{j}(k)\Big{)}+\sqrt{\frac{\log(n/\delta)}{2m}},\quad\forall k\leq\min\bigg{\{}n,\,K,\,\bigg{(}\frac{C}{5\Delta}\bigg{)}^{2}\bigg{\}}$ (B.5)
@@
$\mathbb{P}\Big{(}\bar{y}_{j}\not\in\mathcal{I}^{\mathsf{syn}}_{j}(k)\Big{)}=\mathbb{P}\left(|\bar{y}_{j}-\bar{y}^{\mathsf{syn}}_{j,k}|>\frac{C}{\sqrt{k}}\right).$
Since $\bar{y}_{j}\sim N(\mu_{j},1/n)$ and $\bar{y}^{\mathsf{syn}}_{j,k}\sim N(\mu^{\mathsf{syn}}_{j},1/k)$, then
$\bar{y}_{j}-\bar{y}^{\mathsf{syn}}_{j,k}\sim N\left(\mu_{j}-\mu^{\mathsf{syn}}_{j},~{}\frac{1}{k}+\frac{1}{n}\right).$
@@
$\mathbb{P}\Big{(}\bar{y}_{j}\not\in\mathcal{I}^{\mathsf{syn}}_{j}(k)\Big{)}$ $\leq\mathbb{P}\left(|(\bar{y}_{j}-\bar{y}^{\mathsf{syn}}_{j,k})-(\mu_{j}-\mu^{\mathsf{syn}}_{j})|+\Delta>\frac{C}{\sqrt{k}}\right)$
$=2\Phi\left(-\frac{C/\sqrt{k}-\Delta}{\sqrt{k^{-1}+n^{-1}}}\right)\leq 2\Phi\bigg{(}-\frac{4C/(5\sqrt{k})}{\sqrt{k^{-1}+k^{-1}}}\bigg{)}$ ($\Delta\leq C/(5\sqrt{k})$)
@@
When $m>\log(n/\delta)/(2\xi^{2})$, substituting (B.6) into (B.5) yields that for all $k\leq\min\{n,\,K,\,(\frac{C}{5\Delta})^{2}\}$,
@@
When this happens, we have $\widehat{k}\geq\min\{n,\,K,\,(\frac{C}{5\Delta})^{2}\}$.
@@
$=1-\alpha-\sqrt{\frac{\pi}{8}}\cdot\gamma^{-1}\sqrt{\frac{1}{m}}$
@@
$=\gamma\cdot\mathbb{P}\Big{(}\theta(\psi)\not\in{\mathcal{S}}^{\mathsf{syn}}(k)\Big{)}.$ (B.10)
Therefore, when the event in (B.9) happens,
@@ C.1 The OpinionQA Dataset
The original dataset is categorized into topics such as health, crime/security, and political issues. Ideally, we would want to consider questions from the same category to ensure that they are similar enough. However, the category with the most questions has fewer than 200 questions. We thus consider pooling all questions. The dataset has 1,442 survey questions in total, which is too large for our computational resources. We selected a subset of questions as follows. First, while the number of choices ranges from 2 to 19, most questions have 5 choices. To give a fair comparison and for simplicity, we only consider questions with 5 choices. Second, not all questions have choices that can be clearly ordered in sentiments, such as the following one:
@@
We asked GPT-4o to determine if a question’s choices can be ordered in sentiments and we keep those that have GPT-4o’s affirmative answer. This leaves us with 546 questions. To compensate for the loss of similarity by pooling questions across various topics and to further reduce our computational cost, we selected 400 questions that are “most similar” to each other by embedding the question statements using OpenAI’s text-embedding-3-small, calculating the mean, and selecting the 400 questions with the smallest Euclidean distance to the mean. Out of these 400 questions, 15 questions have various issues with their choices by manual inspection, so we exclude them. This leaves us with 385 questions. All these questions happen to have at least 400 responses.
@@
Table 1: Categories of surveyees’ features in the OpinionQA dataset.
@@
In our experiments, LLMs usually directly gave answers in the required format, e.g., ‘[[2]]’.
@@ C.3 Visualization of Simulation Sample Size Selection
Figure 5: Visualization of simulation sample size selection for the OpinionQA dataset. Here $\alpha=0.1$, $\gamma=0.5$, $K=100$, $c=\sqrt{2}$ and $M=2$. The red dotted horizontal line is threshold $\gamma\alpha=0.05$. The blue curve represents $\{L(k)\}_{k=1}^{K}$. Our approach chooses $\widehat{k}$ as the point at which the blue line first up-crosses the red threshold: $\widehat{k}=57$.
@@ D.1 Miscoverage Probability Proxy
Table 5: Means and standard errors of $\widetilde{L}(\widehat{k})$ over 100 random splits for various LLMs and $\alpha$ over the OpinionQA dataset.
@@
Table 6: Means and standard errors of $\widetilde{L}(\widehat{k})$ over 100 random splits for various LLMs and $\alpha$ over the EEDI dataset.
@@ D.3 Half-Width of Confidence Interval
In Figure 6, Table 9, Figure 7 and Table 10, we present the detailed experiment results for the half-width of the synthetic confidence interval ${\mathcal{S}}^{\mathsf{syn}}(\widehat{k})$.
Figure 6: Confidence interval half-width for various LLMs and $\alpha$ over the OpinionQA dataset.
Table 9: Average half-width (with 95% margin of error in parentheses) of the synthetic confidence interval for various LLMs and $\alpha$ over the OpinionQA dataset. GPT-4o has the shortest confidence interval on average.
@@
Figure 7: Confidence interval half-width for various LLMs and $\alpha$ over the EEDI dataset.
Table 10: Average half-width (with 95% margin of error in parentheses) of the synthetic confidence interval for various LLMs and $\alpha$ over the EEDI dataset. DeepSeek-V3 has the shortest confidence interval on average.